OpenStudy (anonymous):

Find the value of y so that the line passing through (6, y) and (7, 5) has a slope of -3.

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

this is what i have so far: 5-y / 7 - 6 = 4/1 = 4

OpenStudy (anonymous):

idk what to do after that

OpenStudy (anonymous):

here you need first the equation of line you have the slope and and one point

OpenStudy (anonymous):

y-5=-3(x-7)

OpenStudy (anonymous):

so it would be y - 5 = -3x + 21 ?

OpenStudy (anonymous):

y = 26 - 3x ?

OpenStudy (anonymous):

you didnt really help me much but...i got the answer myself -.-

OpenStudy (anonymous):

thansk for the start

OpenStudy (anonymous):

:)) good for you